# Simulating non-Markovian open quantum dynamics by exploiting physics-informed neural network

### 09 Mar 2026 v1

@@ I Introduction
Many-body open quantum systems (OQSs) have attracted wide attention due to their profound applications in diversified fields, including physics, chemistry, materials science, and life sciences. These applications cover various fields such as coherent energy transfer in biological photosystems Engel et al. (2007); Scholes (2010); Hildner et al. (2013), charge transfer in molecular aggregates Kasha (1963); Hestand and Spano (2018), electron transport in single molecular junctions Thoss and Evers (2018); Uzma et al. (2021), multidimensional coherent spectroscopy of condensed phase materials Grégoire et al. (2017); Bruder et al. (2018), correlated quantum matter for quantum information and computation Monroe (2002); Head-Marsden et al. (2020), and precise measurement and control of local spin states in surface-adsorbed molecules Piquero-Zulaica et al. (2022). Accurate characterization of environmental dissipations and system correlations is crucial for the studies of OQSs Wilson (1975); Madhavan et al. (1998); Zhao et al. (2005); Li et al. (2020); Zheng et al. (2013); Yang et al. (2019); Veldman et al. (2021); Ding et al. (2025).
@@
Numerous approaches have been developed to simulate the non-Markovian dynamics of many-body OQSs Tamascelli et al. (2018); Lambert et al. (2019); Strunz et al. (1999); Shao (2004); Han et al. (2019); Stockburger and Grabert (2002); Suess et al. (2014); Cohen et al. (2015); Erpenbeck et al. (2023); Güttge et al. (2013); Cazalilla and Marston (2002); Lode et al. (2020); Wang and Thoss (2009); Weimer et al. (2021); Li et al. (2023); Yan (2014); Li et al. (2024b); Moix and Cao (2013); Duan et al. (2017); Hsieh and Cao (2018); Wang and Thoss (2003). Among these, the dissipaton-embedded quantum master equation (DQME) Yan (2014); Li et al. (2024b) has emerged as a promising method, offering a compact representation of quantum dissipative dynamics and excellent transferability. Its second-quantized formulation has been demonstrated to be equivalent to the extensively employed hierarchical equations of motion (HEOM) Tanimura and Kubo (1989); Yan et al. (2004); Jin et al. (2008); Li et al. (2012), and the dissipaton picture shares common features with the pseudomode theory Garraway (1997); Tamascelli et al. (2018); Lambert et al. (2019); Cirio et al. (2023); Lin et al. (2025); Pleasance et al. (2020). However, these methods all confront an inherent exponential wall: their computational costs scale exponentially with both the system size and the complexity of environmental memory.
@@
To overcome this problem, one common approach is to employ a variational ansatz to represent the dynamical variables in a quantum dynamical equation, such as the reduced density operator (RDO) and auxialiary density operators (ADOs) in the HEOM method and the reduced density tensor (RDT) in the DQME method. Tensor network states (TNSs) Verstraete et al. (2008); Schollwöck (2011); Orús (2014); Shi et al. (2018) and neural quantum states (NQSs) Carleo and Troyer (2017); Schmitt and Heyl (2020); Hartmann and Carleo (2019); Nagy and Savona (2019); Vicentini et al. (2019); Yoshioka and Hamazaki (2019); Cao et al. (2025); Ye et al. (2025) approaches have been developed to efficiently represent these time-dependent dynamical variables. For instance, matrix product states (MPSs) have been employed to compress the ADOs in HEOM Shi et al. (2018); Ke et al. (2022); Ke (2023); Preston et al. (2025), while restricted boltzmann mechines (RBM) have been used to represent the RDT in DQME Cao et al. (2025). The resulting MPS-HEOM and RBM-DQME approaches have achieved significant reduction in the number of required dynamical variables.
@@
When simulating the dynamics of OQSs with variational methods, the time-dependent variational principle (TDVP) Carleo and Troyer (2017); Schmitt and Heyl (2020); Hartmann and Carleo (2019); Reh et al. (2021) is employed to project the reduced system dynamics onto the parameter space of a chosen variational ansatz. This projection yields equations of motion for the variational parameters. A key computational burden in this approach is the necessity to compute the time derivatives of these variational parameters at each incremental time step, which which significantly hinders the broader development and application of variational methods Cao et al. (2025). For instance, applying TDVP within the NQS framework requires solving a large system of linear equations at each time step. Both the construction of the coefficient matrix and the subsequent solution of the system of linear equations are time-consuming Cao et al. (2025) due to the high Monte Carlo sampling number for high-accuracy requirements and the iterative optimizer if the linear problem is large.
@@
An alternative method is provided by physics-informed neural networks (PINNs), which use neural networks to represent solutions of partial differential equations (PDEs) by reformulating the governing equations as a loss function optimization problem. This is achieved by by embedding the mathematical structure of the PDE (e.g., through differential operators such as $\partial_{t},\partial_{x}$) directly into the neural network, and defining a loss that incorporates residuals from the governing equations and boundary conditions. For time-dependent PDEs, time $t$ is treated as an additional input node, allowing a single network to represent the solution over the entire temporal domain. Since the original PINN formulation Raissi (2018), numerous variants have been developed, including physics-constrained neural networks Liu and Wang (2021), hp-variational PINNs Kharazmi et al. (2021), and conservative PINNs Jagtap et al. (2020). PINNs have found broad application across fields such as fluid dynamics, optics, electromagnetism, and geoscience Cuomo et al. (2022). However, their use in quantum mechanics remains largely limited to simple cases such as one-dimensional Schrödinger equations Raissi (2018); Raissi et al. (2019). While recent attempts have been made to apply PINNs to quantum master equations for OQSs Dugan et al. (2023), the results have shown rather limited success. Nevertheless, the PINN strategy offers a potential computational advantage: by encoding temporal evolution directly into the network inputs, it avoids the need to update network parameters at each time step, which is required in NQS approaches that rely on the TDVP. This feature could circumvent the key computational bottlenecks associated with NQS methods.
@@
In this work, we propose a PINN-based framework that represents the dynamics of OQSs governed by DQME via time-independent neural networs, aiming at avoiding the time-consuming TDVP in conventional variational methods. To the best of our konwledge, the study represents the first demonstration of applying PINNs to the dynamics of many-body OQSs, opening a promising avenue for effcient and scalable quantum dynamics simulation.
@@ II.1 Fermionic DQME formalism
We consider an Anderson impurity model Anderson (1961) in which the impurity is linearly coupled to the environment consisting of several noninteracting electron reservoirs. The Hamiltonian of the total system is described by (with $e=\hbar=1$)
@@
Since a reservoir of noninteracting fermionic particles linearly coupled to the system follows Gaussian statistics, its influence on the reduced system dynamics is entirely captured by the reservoir hybridization correlation functions Jin et al. (2008), which are defined as
@@
To be concise, we define the $j$ index as the multi-index $\{\alpha\nu sp\}$ here. We adopt the Padé spectral decomposition scheme Hu et al. (2010) to perform the exponential expansion in Eq. (8).
@@
In the DQME framework, the expansion of the reservoir correlation functions defines the dissipatons (quasiparticles of dissipative effects with complex energies) Hu et al. (2010); Yan (2014); Li et al. (2024b). $j$ labels the dissipaton levels, and $\sigma$ denotes the dissipaton charge ($\sigma=+$ for hole-type; $\sigma=$- for electron-type). The imaginary and real parts of $\gamma_{j}^{\sigma}$ correspond to the dissipaton energy and inverse lifetime, respectively, while $\eta^{\sigma}_{j}$ quantifies the system-dissipaton coupling strength.
Figure 1 illustrates the fermionic DQME theory, where the original OQS is mapped to a dissipaton-embedded system described in terms of the reduced density tensor (RDT), $\bm{\rho}=\rho(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$. Here, vectors $\vec{n}$ and $\vec{n}^{\prime}$ describe system fermion configurations, while $\vec{m}^{-}$ and $\vec{m}^{+}$ represent electron-type and hole-type dissipaton configurations, respectively. The RDT elements with ($\vec{m}^{-},\vec{m}^{+})=\vec{0}$ yield the system’s reduced density operator (RDO). As shown in Fig. 1, each level can accommodate at most one fermion or dissipaton in accordance with the Pauli exclusion principle. In practice, a maximal dissipaton occupation number $M_{\rm max}$ is introduce to facilitate the numerical solution of the DQME Li et al. (2024b).
The fermionic DQME is given by Li et al. (2024b):
@@
Fundamentally, dissipatons share similarities with previously proposed pseudomodes Garraway (1997); Tamascelli et al. (2018); Lambert et al. (2019); Cirio et al. (2023); Lin et al. (2025); Pleasance et al. (2020), and the DQME is formally equivalent to the exact HEOM theory Jin et al. (2008); Li et al. (2012) for reduced system dynamics.
@@ II.2 PINN-DQME method
Figure 2: Schematic of the neural network representing $\rho_{\rm pre}(\vec{n},\vec{n}^{\prime};\vec{m};t)$. The input layer consists of three parts: $\{n_{1},\cdots,n_{N_{{}_{\rm S}}}\}$ and $\{n^{\prime}_{1},\cdots,n^{\prime}_{N_{{}_{\rm S}}}\}$ comprise the system part with $N_{{}_{\rm S}}$ being the number of the system fermion levels, $\{m_{1},\cdots,m_{N_{{}_{\rm E}}}\}$ denote the environmental indeces with $N_{{}_{\rm E}}$ being the number of memory-carrying dissipaton levels, and $\{f_{1}(t),\cdots,f_{N_{{}_{\rm T}}}(t)\}$ denote a group of functions of time. Here, $n_{i}$ and $m_{j}$ represent the occupation number on $i$th system level and $j$th dissipaton level, respectively; which take the value of 0 or 11.
We utilize a multilayer preceptron (MLP) with complex-valued weights and biases to represent the solution of DQME $\rho(\vec{n},\vec{n}^{\prime};\vec{m};t)$, where $\vec{m}=(\vec{m}^{-},\vec{m}^{+})$. The architecture of the MLP is illustrated in Fig. 2, where the input layer consists of the system nodes ($\vec{n},\vec{n}^{\prime}$), the environmental nodes $\vec{m}$ and the time nodes $\vec{f}(t)$. The output is the value of the solution projected onto the basis $(\vec{n},\vec{n}^{\prime};\vec{m})$ at time $t$. Compared to the NQS ansatz, the input layer contains several additional nodes corresponding to the time to encode the information of evolution in the neural network, so this network can represent quantum states at different times rather than at a specific time instant. While the conventional PINN methods Raissi (2018); Cuomo et al. (2022) usually have only one time node $f_{1}(t)=t$ in the input layer, here we use a group of time nodes to improve the capability of the MLP for representing complex time evolution. For an MLP consisting of K1K-1 hidden layers, the computation of its output can be described as a sequence of $K$ matrix-vector multiplications separated by a nonlinear activation function $a:\mathbb{C}\to\mathbb{C}$. Specifically, the RDT of DQME is approximated by
@@
The dimensionality of the RDT can be reduced by leveraging its inherent symmetry and sparsity Hou et al. (2015). First, we enforce the symmetry constraint by constructing
@@
where $\lfloor\cdot\rfloor$ denotes the floor function, $M^{\sigma}=\sum_{j}m_{j}^{\sigma}$, $m_{j}^{\sigma}$ is the $j$-th element of $\vec{m}^{\sigma}$, and $\rho_{\rm pre}(\vec{s};t)$ is the MLP representation of the solutions. Here, $\vec{s}\equiv(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$ represents a visible state, while $\vec{s}^{\,T}\equiv(\vec{n}^{\prime},\vec{n};\vec{m}^{+},\vec{m}^{-})$ is its block-swapped counterpart. Next, we introduce a filter function $\mathcal{F}_{\rm{spa}}(\vec{s})$ that identifies, a priori, the zero elements of $\bm{\rho}$ based on the sparsity pattern dictated by the forms of $H_{{}_{\rm S}}$ and $H_{{}_{\rm SE}}$ Hou et al. (2015), which keeps invariant during the evolution. In practice, $\mathcal{F}_{\rm{spa}}(\vec{s})$ normally reduces the $\vec{s}$-space by over an order of magnitude. The resulting symmetrized and filtered RDT is given by $\rho(\vec{s};t)=\mathcal{F}_{\rm spa}(\vec{s})\rho_{\rm sym}(\vec{s};t)$ Cao et al. (2025).
@@
where $L_{\rm R}$ is the residual term from DQME, $L_{\rm I}$ comes from the initial condition $\bm{\rho}_{\theta}(\tau_{0})=\bm{\rho}(\tau_{0})$, $L_{\rm I}$ is introduced to enforce the physical constraints that the trace of RDO equals 1, and $\omega_{\rm R}$, $\omega_{\rm I}$ and $\omega_{\rm tr}$ are the weights assigned to the equation residual, the initial condition, and the trace residual, respectively. In practice, we empirically choose $\omega_{\rm R}=0.2$, $\omega_{\rm I}=0.8$ and $\omega_{\rm tr}=20$, where not otherwise specified. In these terms, $N_{\rm R}$ is the count of residual points (indexed by $\tau_{i}$), $\bm{\rho}_{\theta}(t)$ is the parametric RDT at time $t$, $\rho_{0}(t)$ is the parametric RDO of the system at time $t$, and ${\rm tr}_{{}_{\rm S}}$ means taking the trace of the system subspace. The time derivative of the RDT, $\dot{\bm{\rho}}$, is computed by numerical differentiation with a sufficiently small time difference of $\delta t=1.5\times 10^{-9}\,\Gamma^{-1}$. $\|\cdot\|_{2}$ denotes the 2-norm, while $\|\cdot\|_{M}^{2}$ is defined as: $\|\bm{\rho}\|_{M}^{2}=\sum_{\vec{s}}e^{\lambda M(\vec{s})}|\rho(\vec{s})|^{2}$ with $M(\vec{s})=\sum_{j}m_{j}$ and the empirical exponent $\lambda=-3$. This exponent is introduced to emphasize the importance of elements with low $M$ in RDT, which hold greater relevance to physical observables. Minimizing this loss function, we can find the possible approximation of the solutions of DQME under specific initial conditions.
In cases where the quantum state of the system undergoes significant changes over the studied time scale, the complex dynamics can render the full evolution difficult to simulate with a single neural network Krishnapriyan et al. (2021), due to the rugged loss landscape arising from the governing equations. To address this, we adopt a domain decomposition strategy Jagtap et al. (2020); Krishnapriyan et al. (2021), which divides the entire time interval into several subdomains and approximates the evolution within each using an individual neural network. The loss function for the $p$-th subdomain is defined as:
@@
There are many methods in machine learning to optimize the loss function, such as stochastic gradient descent (SGD) Robbins and Monro (1951), Adam Kingma and Ba (2014), and AdamW Loshchilov and Hutter (2017). The most common optimization method in PINN is limited-memory Broyden-Fletcher-Goldfarb-Shanno algorithm (L-BFGS) Cuomo et al. (2022). However, all of them failed when applied to the PINN-DQME method due to the significant complexity of DQME. Consequently, we chose the standard BFGS algorithm Fletcher (2013) , which is highly accurate but somewhat expensive, to minimize the loss function involved in the PINN-DQME method.

### 29 Apr 2026 v2

@@ I Introduction
Many-body open quantum systems (OQSs) have attracted wide attention due to their profound applications in diversified fields, including physics, chemistry, materials science, and life sciences. These applications cover various fields such as coherent energy transfer in biological photosystems Engel2007 ; doi:10.1021/jz900062f ; doi:10.1126/science.1235820 , charge transfer in molecular aggregates 6395f2ce-ffb5-3279-8695-df26ad69dc2d ; doi:10.1021/acs.chemrev.7b00581 , electron transport in single molecular junctions 10.1063/1.5003306 ; Uzma2021 , multidimensional coherent spectroscopy of condensed phase materials 10.1063/1.4994987 ; Bruder2018 , correlated quantum matter for quantum information and computation Monroe2002 ; head2020quantum , and precise measurement and control of local spin states in surface-adsorbed molecules RevModPhys.94.045008 . Accurate characterization of environmental dissipations and system correlations is crucial for the studies of OQSs wilson1975renormalization ; doi:10.1126/science.280.5363.567 ; doi:10.1126/science.1113449 ; li2020molecular ; PhysRevLett.111.086601 ; doi:10.1126/science.aay6779 ; doi:10.1126/science.abg8223 ; Ding25084114 .
@@
Numerous approaches have been developed to simulate the non-Markovian dynamics of many-body OQSs Tam18030402 ; Lam193721 ; PhysRevLett.82.1801 ; 10.1063/1.1647528 ; PhysRevLett.123.050601 ; PhysRevLett.88.170407 ; suess2014hierarchy ; PhysRevLett.115.266802 ; PhysRevLett.130.186301 ; PhysRevB.87.115115 ; PhysRevLett.88.256403 ; RevModPhys.92.011001 ; 10.1063/1.3173823 ; RevModPhys.93.015008 ; li2023dissipatons ; Yan14054105 ; li2024toward ; moix2013hybrid ; duan2017zero ; hsieh2018unified ; 10.1063/1.1580111 . Among these, the dissipaton-embedded quantum master equation (DQME) Yan14054105 ; li2024toward has emerged as a promising method, offering a compact representation of quantum dissipative dynamics and excellent transferability. Its second-quantized formulation has been demonstrated to be equivalent to the extensively employed hierarchical equations of motion (HEOM) Tan89101 ; YAN2004216 ; Jin08234703 ; PhysRevLett.109.266403 , and the dissipaton picture shares common features with the pseudomode theory PhysRevA.55.2290 ; Tam18030402 ; Lam193721 ; cirio2023pseudofermion ; Lin251289 ; PhysRevResearch.2.043058 . However, these methods all confront an inherent exponential wall: their computational costs scale exponentially with both the system size and the complexity of environmental memory.
@@
To overcome this problem, one common approach is to employ a variational ansatz to represent the dynamical variables in a quantum dynamical equation, such as the reduced density operator (RDO) and auxiliary density operators (ADOs) in the HEOM method and the reduced density tensor (RDT) in the DQME method. Tensor network states (TNSs) verstraete2008matrix ; schollwock2011density ; orus2014practical ; shi2018efficient and neural quantum states (NQSs) doi:10.1126/science.aag2302 ; PhysRevLett.125.100503 ; PhysRevLett.122.250502 ; PhysRevLett.122.250501 ; PhysRevLett.122.250503 ; PhysRevB.99.214306 ; cao2026simulating ; ye2025simulating approaches have been developed to efficiently represent these time-dependent dynamical variables. For instance, matrix product states (MPSs) have been employed to compress the ADOs in HEOM shi2018efficient ; ke2022hierarchical ; ke2023tree ; preston2025nonadiabatic , while restricted Boltzmann mechines (RBMs) have been used to represent the RDT in DQME cao2026simulating . The resulting MPS-HEOM and RBM-DQME approaches have achieved significant reduction in the number of required dynamical variables. Recently, operator learning techniques have also been explored, offering an alternative approach for simulating OQS dynamics by utilizing machine learning zhang2024artificial ; zhang2025neural .
@@
When simulating the dynamics of OQSs with variational methods, the time-dependent variational principle (TDVP) doi:10.1126/science.aag2302 ; PhysRevLett.125.100503 ; PhysRevLett.122.250502 ; reh2021time is employed to project the reduced system dynamics onto the parameter space of a chosen variational ansatz. This projection yields equations of motion for the variational parameters. A key computational burden in this approach is the necessity to compute the time derivatives of these variational parameters at each incremental time step, which significantly hinders the broader development and application of variational methods cao2026simulating . For instance, applying TDVP within the NQS framework requires solving a large system of linear equations at each time step. Both the construction of the coefficient matrix and the subsequent solution of the system of linear equations are time-consuming cao2026simulating due to the high Monte Carlo sampling number for high-accuracy requirements and the iterative optimizer if the linear problem is large.
@@
An alternative method is provided by physics-informed neural networks (PINNs), which use neural networks to represent solutions of partial differential equations (PDEs) by reformulating the governing equations as a loss function optimization problem. This is achieved by embedding the mathematical structure of the PDE (e.g., through differential operators such as $\partial_{t},\partial_{x}$) directly into the neural network, and defining a loss that incorporates residuals from the governing equations and boundary conditions. For time-dependent PDEs, time $t$ is treated as an additional input node, allowing a single network to represent the solution over the entire temporal domain. Since the original PINN formulation raissi2018deep , numerous variants have been developed, including physics-constrained neural networks liu2021dual , hp-variational PINNs kharazmi2021hp , and conservative PINNs jagtap2020conservative . PINNs have found broad application across fields such as fluid dynamics, optics, electromagnetism, and geoscience cuomo2022scientific . However, their use in quantum mechanics remains largely limited to non-dissipative situations, such as solving the Schrödinger equation for the transverse field Ising model, sinibaldi2026time ; van2025many or addressing data-driven problems ullah2024physics ; ullah2025machine ; atif2026toward . While recent attempts have been made to apply PINNs to quantum master equations for OQSs dugan2023q , the results have shown rather limited success. Nevertheless, the PINN approach offers a potential computational advantage: by encoding temporal evolution directly into the network inputs, it avoids the need to update network parameters at each time step, which is required for NQS approaches that rely on the TDVP. This feature could circumvent the key computational bottlenecks associated with NQS methods.
@@
In this work, we propose a PINN-based framework that represents the dynamics of OQSs governed by DQME via time-independent neural networks, aiming at avoiding the time-consuming TDVP in conventional variational methods. To the best of our knowledge, the study represents the first demonstration of applying PINNs to the dissipative dynamics of many-body OQSs, opening a promising avenue for efficient and scalable quantum dynamics simulation.
@@ II.1 Fermionic DQME formalism
We consider an Anderson impurity model And6141 in which the impurity is linearly coupled to the environment consisting of several noninteracting electron reservoirs. The Hamiltonian of the total system is described by (with $e=\hbar=1$)
@@
Since a reservoir of noninteracting fermionic particles linearly coupled to the system follows Gaussian statistics, its influence on the reduced system dynamics is entirely captured by the reservoir hybridization correlation functions Jin08234703 , which are defined as
@@
To be concise, we define the $j$ index as the multi-index $\{\alpha\nu sp\}$ here. We adopt the Padé spectral decomposition scheme hu2010communication to perform the exponential expansion in Eq. (8).
@@
In the DQME framework, the expansion of the reservoir correlation functions defines the dissipatons (quasiparticles of dissipative effects with complex energies) hu2010communication ; Yan14054105 ; li2024toward . $j$ labels the dissipaton levels, and $\sigma$ denotes the dissipaton charge ($\sigma=+$ for hole-type; $\sigma=$- for electron-type). The imaginary and real parts of $\gamma_{j}^{\sigma}$ correspond to the dissipaton energy and inverse lifetime, respectively, while $\eta^{\sigma}_{j}$ quantifies the system-dissipaton coupling strength.
Figure 1 illustrates the fermionic DQME theory, where the original OQS is mapped to a dissipaton-embedded system described in terms of the reduced density tensor (RDT), $\bm{\rho}=\rho(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$. Here, vectors $\vec{n}$ and $\vec{n}^{\prime}$ describe system fermion configurations, while $\vec{m}^{-}$ and $\vec{m}^{+}$ represent electron-type and hole-type dissipaton configurations, respectively. The RDT elements with ($\vec{m}^{-},\vec{m}^{+})=\vec{0}$ yield the system’s reduced density operator (RDO). As shown in Fig. 1, each level can accommodate at most one fermion or dissipaton in accordance with the Pauli exclusion principle. In practice, a maximal dissipaton occupation number $M_{\rm max}$ is introduce to facilitate the numerical solution of the DQME li2024toward .
The fermionic DQME is given by li2024toward :
@@
Fundamentally, dissipatons share similarities with previously proposed pseudomodes PhysRevA.55.2290 ; Tam18030402 ; Lam193721 ; cirio2023pseudofermion ; Lin251289 ; PhysRevResearch.2.043058 , and the DQME is formally equivalent to the exact HEOM theory Jin08234703 ; PhysRevLett.109.266403 for reduced system dynamics.
@@ II.2 PINN-DQME method
Figure 2: Schematic of the neural network representing $\rho_{\rm pre}(\vec{n},\vec{n}^{\prime};\vec{m};t)$. The input layer consists of three parts: the fermion configurations $\{n_{1},\cdots,n_{N_{{}_{\rm S}}}\}$ and $\{n^{\prime}_{1},\cdots,n^{\prime}_{N_{{}_{\rm S}}}\}$ from the $N_{{}_{\rm S}}$ system levels, the dissipaton configurations $\{m_{1},\cdots,m_{N_{{}_{\rm E}}}\}$ from the $N_{{}_{\rm E}}$ memory-carrying dissipaton levels, and a set of time-varying functions $\{f_{1}(t),\cdots,f_{N_{{}_{\rm T}}}(t)\}$. Here, $n_{i}$ and $m_{j}$ represent the occupation number on $i$th system level and $j$th dissipaton level, respectively; which take the value of 0 or 11. The arrows indicate full connections between layers.
We utilize a multilayer perceptron (MLP) with complex-valued weights and biases to represent the solution of DQME $\rho(\vec{n},\vec{n}^{\prime};\vec{m};t)$, where $\vec{m}=(\vec{m}^{-},\vec{m}^{+})$. The architecture of the MLP is illustrated in Fig. 2, where the input layer consists of the system nodes ($\vec{n},\vec{n}^{\prime}$), the environmental nodes $\vec{m}$ and the time nodes $\vec{f}(t)$. The output is the value of the solution projected onto the basis $(\vec{n},\vec{n}^{\prime};\vec{m})$ at time $t$. Compared to the NQS ansatz, the input layer contains several additional nodes corresponding to the time to encode the information of evolution in the neural network, so this network can represent quantum states at different times rather than at a specific time instant. While the conventional PINN methods raissi2018deep ; cuomo2022scientific usually have only one time node $f_{1}(t)=t$ in the input layer, here we use a group of time nodes to improve the capability of the MLP for representing complex time evolution. For an MLP consisting of K1K-1 hidden layers, the computation of its output can be described as a sequence of $K$ matrix-vector multiplications separated by a nonlinear activation function $a:\mathbb{C}\to\mathbb{C}$. Specifically, the RDT of DQME is approximated by
@@
The dimensionality of the RDT can be reduced by leveraging its inherent symmetry and sparsity Hou15104112 . First, we enforce the symmetry constraint by constructing
@@
where $\lfloor\cdot\rfloor$ denotes the floor function, $M^{\sigma}=\sum_{j}m_{j}^{\sigma}$, $m_{j}^{\sigma}$ is the $j$-th element of $\vec{m}^{\sigma}$, and $\rho_{\rm pre}(\vec{s};t)$ is the MLP representation of the solutions. Here, $\vec{s}\equiv(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$ represents a visible state, while $\vec{s}^{\,T}\equiv(\vec{n}^{\prime},\vec{n};\vec{m}^{+},\vec{m}^{-})$ is its block-swapped counterpart. Next, we introduce a filter function $\mathcal{F}_{\rm{spa}}(\vec{s})$ that identifies, a priori, the zero elements of $\bm{\rho}$ based on the sparsity pattern dictated by the forms of $H_{{}_{\rm S}}$ and $H_{{}_{\rm SE}}$ Hou15104112 , which keeps invariant during the evolution. In practice, $\mathcal{F}_{\rm{spa}}(\vec{s})$ normally reduces the $\vec{s}$-space by over an order of magnitude. The resulting symmetrized and filtered RDT is given by $\rho(\vec{s};t)=\mathcal{F}_{\rm spa}(\vec{s})\rho_{\rm sym}(\vec{s};t)$ cao2026simulating .
@@
where $L_{\rm R}$ is the residual term from DQME, $L_{\rm I}$ comes from the initial condition $\bm{\rho}_{\theta}(\tau_{0})=\bm{\rho}(\tau_{0})$, $L_{\rm{\rm tr}}$ is introduced to enforce the physical constraints that the trace of RDO equals 1, and $\omega_{\rm R}$, $\omega_{\rm I}$ and $\omega_{\rm tr}$ are the weights assigned to the equation residual, the initial condition, and the trace residual, respectively. In practice, we empirically choose $\omega_{\rm R}=0.2$, $\omega_{\rm I}=0.8$ and $\omega_{\rm tr}=20$, where not otherwise specified. In these terms, $N_{\rm R}$ is the count of residual points (indexed by $\tau_{i}$), $\bm{\rho}_{\theta}(t)$ is the parametric RDT at time $t$, $\rho_{0}(t)$ is the parametric RDO of the system at time $t$, and ${\rm tr}_{{}_{\rm S}}$ means taking the trace of the system subspace. The time derivative of the RDT, $\dot{\bm{\rho}}$, is computed by numerical differentiation with a sufficiently small time difference of $\delta t=1.5\times 10^{-9}\,\Gamma^{-1}$. $\|\cdot\|_{2}$ denotes the 2-norm, while $\|\cdot\|_{M}^{2}$ is defined as: $\|\bm{\rho}\|_{M}^{2}=\sum_{\vec{s}}e^{\lambda M(\vec{s})}|\rho(\vec{s})|^{2}$ with $M(\vec{s})=\sum_{j}m_{j}$ and the empirical exponent $\lambda=-3$. This exponent is introduced to emphasize the importance of elements with low $M$ in RDT, which hold greater relevance to physical observables. Minimizing this loss function, we can find the possible approximation of the solutions of DQME under specific initial conditions.
In cases where the quantum state of the system undergoes significant changes over the studied time scale, the complex dynamics can render the full evolution difficult to simulate with a single neural network krishnapriyan2021characterizing , due to the rugged loss landscape arising from the governing equations. To address this, we adopt a domain decomposition strategy jagtap2020conservative ; krishnapriyan2021characterizing , which divides the entire time interval into several subdomains and approximates the evolution within each using an individual neural network. The loss function for the $p$-th subdomain is defined as:
@@
There are many methods in machine learning to optimize the loss function, such as stochastic gradient descent (SGD) robbins1951stochastic , Adam kingma2014adam , and AdamW loshchilov2017decoupled . The most common optimization method in PINN is limited-memory Broyden-Fletcher-Goldfarb-Shanno algorithm (L-BFGS) cuomo2022scientific . However, all of them failed when applied to the PINN-DQME method due to the significant complexity of DQME. Consequently, we chose the standard BFGS algorithm fletcher2013practical , which is highly accurate but somewhat expensive, to minimize the loss function involved in the PINN-DQME method.